# Cheick Tidiane Ba, Richard G. Clegg

### v1

@@ 4.1 Dataset
Figure 1: Daily price of the currencies under study, in the periods covered by the full dataset. We represent the closing prices on April 1st and onward, normalising the values of each by dividing by the starting value in the observed period. In the insets, we provide a zoomed version around the value of 1 USD.
4.2 Analysis methods
The data is analysed as a directed, multilayer temporal graph. A node represents an Ethereum wallet. A directed edge represents a transaction between two wallets, the weight represents the amount of currency, the direction represents the direction money is transferred and the layer represents the individual currency. Using the Raphtory system [37] we can easily perform analysis using sliding windows to consider the trades within different parts of the data and investigate how the temporal graph behaves within a time period [^3] Code will be made publicly available in case of acceptance.. If we define a time window $T=\left[t_{1},t_{2}\right)$ then that window will include trades made at all times $t\leq t_{1}$ and $t<t_{2}$. [^4] The closed/open interval means we do not double count trades that occur exactly on a border if, for example, we move a window forward every hour we do not want trades occurring exactly on the hour to be in two different windows. The degree of a node within a time window is a measure of the number of trading partners that the wallet has within that time window. The balance of the weights of inbound minus outbound link weights is the node’s income during the period.
@@ 4.2 Analysis methods
Figure 2: Illustration of the construction of the temporal multilayer network. Given a set of transactions, we can model them as a network, where each layer represents one of the currencies under study.
@@
For analysis purposes, we want to separate the data into a pre-crash and post-crash period to investigate the normal behaviour of the market and to see if it returns to the same situation after the crash occurs. Obviously, it is somewhat arbitrary as to when such an event ends. For the purposes of this paper, we have chosen the 9th May 2022 as the centre of the crash event and we exclude that day and a week on either side. This leads to a fifteen-day exclusion period with the pre-crash period including all data up to the end of 1st May and the post-crash period including all data from the beginning of the 17th May. This gives us thirty days of complete data before the exclusion zone and by considering all data up to the end of 15th June we have thirty days after the exclusion zone. The exclusion zone is plotted on all graphs as a light red area with $C$ marking the centre. Graphs are also marked with $S1$ and $S2$ for the two major selling events on 3rd April 2022 and 19th April 2022 that preceded the crash (see the timeline in Section 2) and with $T2$ for the launch of Terra 2.0 blockchain on 27th May 2022.
4.3 Basic network statistics
@@
Table 2: Network statistics in the full graph and each layer. Active: source (out) or target (in) of at least one transaction. Sinks perform no out action, while sources perform only out actions
@@
Figure 3: Probability distribution for node degree on the full data set, pre-crash period and post-crash period. The axes are logarithmic on both scales.
@@ 5 Results
5.1 What is the nature of the cryptocurrency ecosystem before and after the attack
@@
Table 3: Active users. We represent in how many layers are users active.
@@
Figure 4: Cross-correlation $\max_{k}(\rho(k))$ of trade patterns for the six currencies studied.
@@
From the analysis of Figure 4 the most obvious conclusion is that many of the currencies have an extremely high correlation in the pre-crash period. The number of trades and graph edges (pairs of trading partners) show extremely high correlation in this period. The graph of total trade value shows less correlation even in the pre-crash period but the currency USDT is a marked anomaly having low correlation compared with the others. The likely reason for this is a small number of large volume trade events in the pre-crash period, that appears to be part of the attempt to destabilise the WLUNC, USDT system. In the post crash period, the system has split into two with the remaining four stablecoins extremely highly correlated with each other, but USDT and WLUNC have a much lower correlation. It is also interesting to note that WLUNC and USDT correlate well with each other in almost all analyses apart from the pre-crash trade value which is disrupted for USDT as previously described. The major conclusion of this analysis is that pre-crash all currencies exhibited extremely similar trade patterns (by this measure at least) but post crash WLUNC and USDT have become a separate eco-system.
5.2 Can we anticipate the crash by looking at the cryptocurrencies in the days leading up to it?
@@
Figure 5: Fluctuations in trading volume where each line represents one of the currencies under study. The number of transactions is the total transactions that day. The number of edges is the number of unique pairs of buyers transacting. The tokens transferred are the raw number of tokens and the value in USD uses the close price of that day’s trading to convert this.
@@
5.3 What are the effects of the attack on the system beyond the most directly affected currencies?
@@
Figure 6: Monitoring the new edges between unique user pairs across different window sizes. Here new means that this is the first window in which they are present. On the X-axis: Time. On the Y-axis: number of edges according to the selected time window.
Our objective is to examine the driving forces behind the system, discerning whether it is propelled by returning users who consistently contribute or by new users who have not been encountered previously. To achieve this, we adopt a methodology inspired by prior research  [38], wherein distinctions are made between new users or edges within a specified window. These entities encompass users and pairwise interactions occurring within the window and have not been active at any time prior to the selected window. The adjustment of the window size $\tau$ enables varied definitions of an active user. We monitor, for each layer, the number of new interactions according to daily, weekly, and bi-weekly windows. An edge is considered new if it is the first time window it occurs in.
We show the main currencies WLUNC and USTC, in Figures  6 and  6 respectively. In both currencies, an intriguing pattern emerges. There are two peaks in the number of new edges, with the first peak coinciding with the crash period. This is likely a result of currency holders responding by engaging in trades with new users (disposing of the coin) while new users enter the system, likely driven by speculative motives. The second notable surge in USTC and WLUNC plots aligns with the launch of Terra 2.0 discussed in Section 2. A plausible explanation for this second surge stems from the criteria for distributing new LUNA tokens on Terra 2.0 [6, 7]. The proposal specifies eligibility criteria for users to receive tokens, based on selected snapshots pre-attack and post-attack. The Governance chose 7th May 2022 at 11 pm for the pre-crash snapshot and 27th May 2022 at 3.59 AM for the post-crash snapshot. Notably, this choice favours the users who were invested in Terra or Luna before the crash. With the voting concluding on 25th May 2022, making the plan official, and the target launch date on 27th May 2022, users who sold during the crash but intended to join Terra 2.0 had an incentive to purchase WLUNC tokens before the 27th May 2022 deadline. This potentially explains the spikes in USTC and WLUNC on 26th May 2022.
@@
5.4 Can the after-effects of the attack be seen on graph structure as a whole?
@@
Figure 7: Graph measures over time. We monitor different graph measures over time, For each day, the interactions on each currency are treated as an individual graph. On the X-axis: Time. On the Y-axis: the selected graph measures values over time. For clustering coefficient (c), one of the layers exhibits outlier values, so we focus on the main currencies in the main plot, while we provide the complete Figure as an inset.
@@
5.5 Where did user interest move post crash?
@@
(b) Number of layers ((users that use WLUNC)
@@
Figure 8: Number of layers in which users are active. At each timestep, we check the number of layers in which the users are active, giving them a label (1…6). Each line represents the changes in the possible cases (1…6). On the X-axis: Time. On the Y-axis: fraction of each possible value.
@@
Figure 9: Changes of favourite layer for users selected pre-crash and observed post-crash. Each graph is focused on a layer: given a layer, we select users that had at least 1 transaction on that layer. A node is labelled with its favourite layer i.e. the layer in which they have the most Edges.
To gain a deeper understanding of how user behaviour evolved, we conducted an additional analysis focusing on users’ favourite layers—indicating the layer where they are most active—before and after the crash. We presume that users may shift their attention to different currencies post-crash, especially away from the currencies most involved in the crash. The shifts are visualised using Sankey diagrams in Figure 9. The Sankey diagrams reveal that users active before the crash tend to be inactive after the crash. Across all currencies, among those who remain active, there is a notable trend of shifting focus to different currencies; however, there are differences among currencies. Specifically, in the currencies directly involved in the crash, Terra and Luna (see Figures 9 and 9) we can observe how among those who remain active, there is a notable trend of shifting focus to different currencies. For example, in the case of WLUNC, users predominantly migrate to USDC and USDT. Instead, on the other layers (DAI, USDC, USDT, USDP), a higher proportion of users remain faithful, showing less migration to other currencies.
5.6 Can we locate individuals associated with the crash?
@@
Figure 10: Top 10 addresses in terms of USTC tokens sold during selected days. We represent the 2 anomalous days in terms of pre-crash USTC selling, 3rd April 2022 and 19th April 2022, and some control days without anomalies. We represent the percentage of tokens sold over the selected day by the top 10 users, the rest are condensed as the tail. Users from the top 10 that were active on both anomalous days are coloured in red, while users recurrent on other days are coloured in green.
FFigure 10 illustrates the segmentation of the top token sellers on both the anomalous days and the control days. Observations indicate that on peak days (3rd April 2022, 19th April 2022), a few top users dominate the token movement, accounting for a significant percentage of total token sales — almost $99\%$ on both days. In contrast, on control days, sales are more evenly distributed, with the top 10 users contributing a lesser percentage to total sales, we observe a maximum of around $65\%$. This discrepancy suggests that a limited number of users are responsible for the peak sales, showcasing substantial token movements not witnessed on other days.
@@
Another important aspect is the presence of recurrent users across the various days. In general, we observe that some users are active across different days. Users A, B, and C consistently appear in the top 10 during the two anomalous days, driving the anomalous peaks. However, their activities on non-peak days are comparatively subdued. Notably, user C displays activity between peaks but not afterwards, resurfacing only on the last control day. Other accounts intermittently appear in the top 10 on regular days, yet their influence is eclipsed by the substantial movements of users A, B, and C. This reveals that among the users contributing to anomalous peaks, some exhibit coordinated efforts, appearing active exclusively on peak days or displaying interconnected movements.
6 Conclusions
@@
This work underscores how novel temporal network analysis can reveal new details in data from Web3, one of the novel paradigms for the Web. It emphasises the importance of separating transactions in layers in the analysis of temporal heterogeneous data, such as transaction data from the web. Using the Raphtory software we were able to simultaneously look at multiple layers (currencies), and multiple timescales and track events as time evolved.

### v2

@@ 4.1. Dataset
Figure 1. Daily price of the currencies under study, in the periods covered by the full dataset. We represent the closing prices on April 1st and onward, normalising the values of each by dividing by the starting value in the observed period. In the insets, we provide a zoomed version around the value of 1 USD.
4.2. Analysis methods
The data is analysed as a directed, multilayer temporal graph. A node represents an Ethereum wallet. A directed edge represents a transaction between two wallets, the weight represents the amount of currency, the direction represents the direction in which money is transferred and the layer represents the individual currency. This is illustrated in Figure 2.
@@ 4.2. Analysis methods
Figure 2. Illustration of the construction of the temporal multilayer network. Given a set of transactions, we can model them as a network, where each layer represents one of the currencies under study.
Using the Raphtory system (Steer et al., 2023) we can easily perform analysis using sliding windows to consider the trades within different parts of the data and investigate how the temporal graph behaves within a time period [^3] Code will be made publicly available in case of acceptance.. If we define a time window $T=\left[t_{1},t_{2}\right)$ then that window will include trades made at all times $t\leq t_{1}$ and $t<t_{2}$. [^4] The closed/open interval means we do not double count trades that occur exactly on a border if, for example, we move a window forward every hour we do not want trades occurring exactly on the hour to be in two different windows. The degree of a node within a time window is a measure of the number of trading partners that the wallet has within that time window. The balance of the weights of inbound minus outbound link weights is the node’s income during the period.
@@
For analysis purposes, we want to separate the data into a pre-crash and post-crash period to investigate the normal behaviour of the market and to see if it returns to the same situation after the crash occurs. Obviously, when such an event ends is somewhat arbitrary. For the purposes of this paper, we have chosen the 9th May 2022 as the centre of the crash event and we exclude that day and a week on either side. This leads to a fifteen-day exclusion period with the pre-crash period including all data up to the end of 1st May and the post-crash period including all data from the beginning of the 17th May. This gives us thirty days of complete data before the exclusion zone and by considering all data up to the end of 15th June we have thirty days after the exclusion zone. In Figure 3, we present a summary of the methodology and selected filters for reproducibility.
Figure 3. Methodology pipeline. The illustration summarises the selected time filters and the different temporal versions of the graphs we analyse.
The exclusion zone is plotted on all graphs as a light red area with $C$ marking the centre. Graphs are also marked with $S1$ and $S2$ for the two major selling events on 3rd April 2022 and 19th April 2022 that preceded the crash (see the timeline in Section 2) and with $T2$ for the launch of Terra 2.0 blockchain on 27th May 2022.
4.3. Basic network statistics
@@
Table 2. Network statistics in the full graph and each layer. Active: source (out) or target (in) of at least one transaction. Sinks perform no out action, while sources perform only out actions
@@
Figure 4. Probability distribution for node degree on the full data set, pre-crash period and post-crash period. The axes are logarithmic on both scales.
@@ 5. Results
5.1. What is the nature of the cryptocurrency ecosystem before and after the attack
@@
Table 3. Active users. We represent in how many layers are users active.
@@
Figure 5. Cross-correlation $\max_{k}(\rho(k))$ of trade patterns for the six currencies studied.
@@
From the analysis of Figure 5 the most obvious conclusion is that many of the currencies have an extremely high correlation in the pre-crash period. The number of trades and graph edges (pairs of trading partners) show extremely high correlation in this period. Some degree of correlation might be expected since currency fluctuations are driven by the same external events (for example earlier crashes). However, the extremely high correlations in the left hand side of Figure 5(a) and (b) are remarkably high and this correlation can be clearly seen looking at the raw transaction traces on the left hand side of Figure 6(a). The graph of total trade value shows less correlation even in the pre-crash period but the currency USDT is a marked anomaly having low correlation compared with the others. The likely reason for this is a small number of large volume trade events in the pre-crash period, that appears to be part of the attempt to destabilise the WLUNC, USDT system. In the post crash period, the system has split into two with the remaining four stablecoins extremely highly correlated with each other, but USDT and WLUNC have a much lower correlation. It is also interesting to note that WLUNC and USDT correlate well with each other in almost all analyses apart from the pre-crash trade value which is disrupted for USDT as previously described. The major conclusion of this analysis is that pre-crash all currencies exhibited extremely similar trade patterns (by this measure at least) but post crash WLUNC and USDT have become a separate eco-system.
5.2. Can we anticipate the crash by looking at the cryptocurrencies in the days leading up to it?
@@
Figure 6. Fluctuations in trading volume where each line represents one of the currencies under study. The number of transactions is the total transactions that day. The number of edges is the number of unique pairs of buyers transacting. The tokens transferred are the raw number of tokens and the value in USD uses the close price of that day’s trading to convert this.
@@
5.3. What are the effects of the attack on the system beyond the most directly affected currencies?
@@
Figure 7. Monitoring the new edges between unique user pairs across different window sizes. Here new means that this is the first window in which they are present. On the X-axis: Time. On the Y-axis: number of edges according to the selected time window.
Our objective is to examine the driving forces behind the system, discerning whether it is propelled by returning users who consistently contribute or by new users who have not been encountered previously. To achieve this, we adopt a methodology inspired by prior research  (Arnold et al., 2021), wherein distinctions are made between new users or edges within a specified window. These entities encompass users and pairwise interactions occurring within the window and have not been active at any time prior to the selected window. The adjustment of the window size $\tau$ enables varied definitions of an active user. We monitor, for each layer, the number of new interactions according to daily, weekly, and bi-weekly windows. An edge is considered new if it is the first time window it occurs in.
We show the main currencies WLUNC and USTC, in Figures  7 and  7 respectively. In both currencies, an intriguing pattern emerges. There are two peaks in the number of new edges, with the first peak coinciding with the crash period. This is likely a result of currency holders responding by engaging in trades with new users (disposing of the coin) while new users enter the system, likely driven by speculative motives. The second notable surge in USTC and WLUNC plots aligns with the launch of Terra 2.0 discussed in Section 2. A plausible explanation for this second surge stems from the criteria for distributing new LUNA tokens on Terra 2.0 (Terra, 2022; MC, 2022). The proposal specifies eligibility criteria for users to receive tokens, based on selected snapshots pre-attack and post-attack. The Governance chose 7th May 2022 at 11 pm for the pre-crash snapshot and 27th May 2022 at 3.59 AM for the post-crash snapshot. Notably, this choice favours the users who were invested in Terra or Luna before the crash. With the voting concluding on 25th May 2022, making the plan official, and the target launch date on 27th May 2022, users who sold during the crash but intended to join Terra 2.0 had an incentive to purchase WLUNC tokens before the 27th May 2022 deadline. This potentially explains the spikes in USTC and WLUNC on 26th May 2022.
@@
5.4. Can the after-effects of the attack be seen on graph structure as a whole?
@@
Figure 8. Graph measures over time. We monitor different graph measures over time, For each day, the interactions on each currency are treated as an individual graph. On the X-axis: Time. On the Y-axis: the selected graph measures values over time. For clustering coefficient (c), one of the layers exhibits outlier values, so we focus on the main currencies in the main plot, while we provide the complete Figure as an inset.
@@
5.5. Where did user interest move post crash?
@@
(b) Number of layers (users that use WLUNC)
@@
Figure 9. Number of layers in which users are active. At each timestep, we check the number of layers in which the users are active, giving them a label (1…6). Each line represents the changes in the possible cases (1…6). On the X-axis: Time. On the Y-axis: fraction of each possible value.
@@
Figure 10. Changes of favourite layer for users selected pre-crash and observed post-crash. Each graph is focused on a layer: given a layer, we select users that had at least 1 transaction on that layer. A node is labelled with its favourite layer i.e. the layer in which they have the most Edges.
To gain a deeper understanding of how user behaviour evolved, we conducted an additional analysis focusing on users’ favourite layers—indicating the layer where they are most active—before and after the crash. We presume that users may shift their attention to different currencies post-crash, especially away from the currencies most involved in the crash. The shifts are visualised using Sankey diagrams in Figure 10. The Sankey diagrams reveal that users active before the crash tend to be inactive after the crash. Across all currencies, among those who remain active, there is a notable trend of shifting focus to different currencies; however, there are differences among currencies. Specifically, in the currencies directly involved in the crash, Terra and Luna (see Figures 10 and 10) we can observe how among those who remain active, there is a notable trend of shifting focus to different currencies. For example, in the case of WLUNC, users predominantly migrate to USDC and USDT. Instead, on the other layers (DAI, USDC, USDT, USDP), a higher proportion of users remain faithful, showing less migration to other currencies.
5.6. Can we locate individuals associated with the crash?
@@
Figure 11. Top 10 addresses in terms of USTC tokens sold during selected days. We represent the 2 anomalous days in terms of pre-crash USTC selling, 3rd April 2022 and 19th April 2022, and some control days without anomalies. We represent the percentage of tokens sold over the selected day by the top 10 users, the rest are condensed as the tail. Users from the top 10 that were active on both anomalous days are coloured in red, while users recurrent on other days are coloured in green.
@@
Figure 11 illustrates the segmentation of the top token sellers on both the anomalous days and the control days. Observations indicate that on peak days (3rd April 2022, 19th April 2022), a few top users dominate the token movement, accounting for a significant percentage of total token sales — almost $99\%$ on both days. In contrast, on control days, sales are more evenly distributed, with the top 10 users contributing a lesser percentage to total sales, we observe a maximum of around $65\%$. Another important aspect is the presence of recurrent users across the various days. In general, we observe that some users are active across different days. Users A, B, and C consistently appear in the top 10 during the two anomalous days, driving the anomalous peaks. However, their activities on non-peak days are comparatively subdued. Notably, user C displays activity between peaks but not afterwards, resurfacing only on the last control day. Other accounts intermittently appear in the top 10 on regular days, yet their influence is eclipsed by the substantial movements of users A, B, and C. This reveals that among the users contributing to anomalous peaks, some exhibit coordinated efforts, appearing active exclusively on peak days or displaying interconnected movements. Given the extremely unusual distribution of sales (five/six) wallets with approximately the same number of sales dominating $99\%$ of the market there can be little doubt that the individuals (or single individuals or organisation) controlling these wallets is deliberately participating in an attempt to destabilise Luna/Terra. It is also important to remember though that this study is only on those transactions over ethereum and this is discussed in more detail in the next section.
5.7. Discussion
It is important to note the key limitations of this study. The most important one is that we do not directly have access to data from the Luna/Terra ecosystem. The Luna/Terra transactions we study are “wrapped” transactions in WLUNC/USTC traded on the ethereum system and the original data is not available to researchers. While it is impossible to be one hundred percent certain that the answers to our research questions would transfer directly it seems likely that the majority would. There is no economic reason why the trading patterns should be particularly different. With access to the original data then other wallets would presumably be found associated with the attack as discussed in our final research question. It would be very difficult to prove if the wallets we found belong to one individual, one organisation or multiple individuals and it would be even more difficult to do that across different chains even if that data were available.
How would the techniques used here transfer to investigation of other cryptocurrency attacks? It is very hard to say as each attack has a different nature. The attack investigated here was interesting to study precisely because it was attack based upon selling patterns and attacking the algorithmic behaviour of the cryptocurrency associated with the stablecoin. Other cryptocurrency failures have been based on problems which would certainly manifest differently (for example theft from wallets or security problems). We would not usually expect to be able to identify events such as S1 or S2 leading up to such events. However, the broad range of techniques used in this paper will surely prove useful: correlations between currencies, unusual patterns in trading, mapping the transition between currencies as users lose confidence. The ability to do network analysis at different time-scales, across time and between different layers is really the heart of the approach we advocate rather than the exact analyses included in this paper.
6. Conclusions
@@
This work underscores how novel temporal network analysis can reveal new details in data from crytocurrency transactions, one of the novel paradigms for the Web. It emphasises the importance of separating transactions in layers in the analysis of temporal heterogeneous data, such as transaction data from the web. Using the Raphtory software we were able to simultaneously look at multiple layers (currencies), and multiple timescales and track events as time evolved.